40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 24 02:58:46 2011, kiwamu, Update, Green Locking, X arm locked ! time_series.pnggreen_one_arm.png
    Reply  Mon Jan 24 09:33:08 2011, rana, Update, Green Locking, X arm locked ! 
       Reply  Mon Jan 24 13:08:07 2011, kiwamu, Update, Green Locking, Re: X arm locked ! e_pll_oltf.png
    Reply  Mon Jan 24 10:19:21 2011, Koji, Update, Green Locking, X arm locked ! 
       Reply  Mon Jan 24 14:27:13 2011, kiwamu, Update, Green Locking, Re: X arm locked ! masslock_oltf.png
Message ID: 4193     Entry time: Mon Jan 24 10:19:21 2011     In reply to: 4191     Reply to this: 4196
Author: Koji 
Type: Update 
Category: Green Locking 
Subject: X arm locked ! 

Well... The ALS loop is engaged and the error was suppressed.
So, how is the IR error signal stabilized when the IR is brought in to the resonance?

I can see the linear trend of 0.1V/s from 5s to 10s.  This corresponds to 100kHz/s and 13nm
for the residual beat drift and the arm length motion, respectively. That sounds huge. The DC gain must be increased.

ELOG V3.1.3-